OpenStudy (anonymous):

Marcia has $20 to purchase food for a party. Each bag of candy costs $3 and each bottle of soda costs $2. If x is the number of bags of candy she will purchase and y is the number of bottles of soda, which of the following could be a possible solution to the inequality? A. (3, 5) B. (10, 1) C. (1, 10) D. (4, 10)

OpenStudy (anonymous):

3x + 2y <= 20 now sub your answer choices into the inequality and see which one satisfies it. lets try (3,5) x = 3 and y = 5 now sub 3(3) + 2(5) <= 20 9 + 10 <= 20 19 <= 20 --- correct...this one works Now try (10,1) 3(10) + 2(1) <= 20 30 + 2 <= 20 32 <= 20 (incorrect...not this one) Now try (1,10) 3(1) + 2(10) <= 20 3 + 20 <= 20 23 <= 20 (incorrect...not this one) last one...(4,10) 3(4) + 2(10) <= 20 12 + 20 <= 20 32 <= 20 (incorrect...not this one) There appears to be only 1 that will work and that is (3,5) any questions ?